OpenStudy (anonymous):

find inverse of y=(x+3)^2

OpenStudy (anonymous):

@madhu.mukherjee.946

Nnesha (nnesha):

two equal signs ?

Nnesha (nnesha):

(x+3)^2 or (x-3)^2 ??

OpenStudy (anonymous):

(x+3)^2

Nnesha (nnesha):

okay cool same like last one (x+3)^2 is same as (x+3)(x+3) so apply the foil method to expand (x+3)^2 and then replace x with y and y with x

OpenStudy (madhu.mukherjee.946):

@Nnesha pls help me with a stats question i'm stuck i've already tagged u

Nnesha (nnesha):

stat >.>

OpenStudy (anonymous):

so i foiled (x+3)^2 I got x^2+6x+9 so if you switch the x and y you would get x=y^2+6y-9

Nnesha (nnesha):

wait a sec i said something wrong

OpenStudy (dinamix):

\[f(x)^{-1} =\sqrt{x}-3 \]for x>3 and \[f(x)^{-1} =-\sqrt{x}-3\] for x<3

OpenStudy (anonymous):

how do yo get that answer @dinamix

OpenStudy (dinamix):

@sarahm22

OpenStudy (dinamix):

only i use my mind cuz its easy for me this things

OpenStudy (dinamix):

@Nnesha did u see my answer

Nnesha (nnesha):

No.

OpenStudy (anonymous):

what are the steps to getting the answer

OpenStudy (dinamix):

for ops -3 not 3 sorry i type cery fast

Nnesha (nnesha):

how he got the answer: take square root both sides

Nnesha (nnesha):

that's what he did: replace x with y and y with x 2nd) then take square root both sides to cancel out the square

OpenStudy (dinamix):

\[\sqrt{[x]^2} = |x|\]

OpenStudy (dinamix):

did u know this rule right ?

OpenStudy (dinamix):

@Nnesha

Nnesha (nnesha):

yes ?

OpenStudy (dinamix):

\[\sqrt{(x)^2} = |x|\] did u know this rule @Nnesha

Nnesha (nnesha):

yes. I said NO .cuz i didn't know how you got +3

OpenStudy (dinamix):

\[y = (x+3)^2---> \sqrt{y}=|x+3| --> x=\sqrt{y}-3 --> f(x)^{-1} = \sqrt{x}-3\]

OpenStudy (dinamix):

@Nnesha

OpenStudy (dinamix):

this for only when x>-3

OpenStudy (dinamix):

did u see my steps ?

Nnesha (nnesha):

yes \[\huge\rm \sqrt{ x}=\pm \sqrt{(y+3)^2}\] just in case if he isn't familiar with the absolute value

OpenStudy (dinamix):

so my method is wrong @Nnesha o.O ?

Nnesha (nnesha):

i didn't say that :=)

OpenStudy (dinamix):

so what solution @Nnesha i think i am right cuz no mistakes

Nnesha (nnesha):

yes i think so. good job and thanks

OpenStudy (dinamix):

u know how do it for x<-3 right ?@Nnesha

OpenStudy (dinamix):

@Nnesha

Nnesha (nnesha):

yes i do

OpenStudy (dinamix):

@freckles my method is right ? look my answer plz

OpenStudy (dinamix):

its not 3 i forget - okey

OpenStudy (freckles):

I guess he was looking for inverse relation since an inverse function doesn't exist. And yeah... \[y=(x+3)^2 \\ \text{ take square root of both sides } \\ \pm \sqrt{y}=x+3 \\ \text{ now if } x+3 \text{ is positive or zero } (x+3 \ge 0) \\ \text{ then } x+3=\sqrt{y} \\ \text{ and if } x+3 \text{ is negative } \\ \text{ then } x+3=-\sqrt{y} \] \[\text{ solving each } \\ x=\sqrt{y}-3 \text{ if } x+3 \ge 0 \text{ or solving the inequality we get } x \ge -3 \\ x=-\sqrt{y}-3 \text{ if } x+3 <0 \text{ or solving that inequality we get } x<-3 \\ \text{ so we have } y=\sqrt{x}-3 \text{ if } x \ge -3 \\ \text{ or } y=-\sqrt{x}-3 \text{ if } x<-3 \] which you already said

Nnesha (nnesha):

i didn't say tht ur answer is incorrect ....

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @dinamix its not 3 i forget - okey \(\color{blue}{\text{End of Quote}}\) \(\color{blue}{\text{Originally Posted by}}\) @dinamix for ops -3 not 3 sorry i type cery fast \(\color{blue}{\text{End of Quote}}\) i know.

OpenStudy (dinamix):

lol why do that @Nnesha